# K3 surfaces with real or complex multiplication

### v2

@@ 7. Totally real fields
If $E$ is a totally real field, we denote by $\Sigma_{E}$ be the set of real embeddings of $E$; we have $E\otimes_{{\mathbf{Q}}}\mathbf{R}=\underset{\sigma\in\Sigma_{E}}{\prod}E_{\sigma}$, with $E_{\sigma}={\bf R}$ for all $\sigma\in\Sigma_{E}$. If $W$ is a quadratic form over $E$, then $W\otimes_{{\mathbf{Q}}}\mathbf{R}$ decomposes as an orthogonal sum $W\otimes_{{\mathbf{Q}}}\mathbf{R}=\underset{\sigma\in\Sigma_{E}}{\oplus}W_{\sigma}$; each of the $W_{\sigma}$ is a quadratic form over $\mathbf{R}$.
@@
$\bullet$ $md<{\rm dim}(V)-2$,
$\bullet$ $md={\rm dim}(V)-2$, and moreover for $p=2$ and for all prime numbers $p$ dividing ${\rm N}_{E/{\mathbf{Q}}}({\rm det}(W))$, we have ${\rm det}(\mathrm{T}(W))\not=-{\rm det}(V)$ in ${\mathbf{Q}}_{p}^{\times}/{\mathbf{Q}}_{p}^{\times 2}$.
@@
$V\simeq{\mathrm{T}}(W)\oplus V^{\prime}\leavevmode\nobreak\ .$
Moreover $W$ can be chosen in such a way that there is an embedding $\sigma:E\to\mathbf{R}$ with $W_{\sigma}$ of signature $(2,m-2)$.
Proof. Let $X$ be a quadratic form over ${\mathbf{Q}}$ of signature $(1,d-1)$ such that ${\rm det}(X)=1$ and $w(X)=1$ at the prime $2$ and at infinity, and 00 elsewhere; this is possible by Theorem 1.1, (i). By a result of Krüskemper [Kr 92] Theorem A there exists a 1-dimensional quadratic form $Y$ over $E$ of such that $\mathrm{T}(Y)\simeq X$; indeed, the hypotheses of this result are satisfied, since $w(X)=0$ at all non-dyadic primes. Set $U_{1}=X\oplus X$ and $W_{1}=Y\oplus Y$; we have $\mathrm{T}(W_{1})\simeq U_{1}$. The quadratic form $U_{1}$ has signature $(2,2d-2)$, trivial determinant and Hasse invariant.
[…]
Set $W=W_{1}\oplus W_{2}$ and $U=U_{1}\oplus U_{2}$. Note that $U$ is a quadratic form over ${\mathbf{Q}}$ of signature $(2,19)$ with ${\rm det}(U)=-h$ and $w(U)=w(V)$. We have $V^{\prime}=\langle-{\rm det}(U)\rangle$; we claim that $V\simeq U\oplus V^{\prime}$. Indeed, these forms have the same dimension, determinant and signature. We have $w(U\oplus V^{\prime})=w(U)+w(V^{\prime})+({\rm det}(U),{\rm det}(V^{\prime}))$. Since ${\rm det}(V^{\prime})=-{\rm det}(U)$, we have $({\rm det}(U),{\rm det}(V^{\prime})=0$. Moreover, ${\rm dim}(V^{\prime})=1$, hence $w(V^{\prime})=0$. This implies that $w(U\oplus V^{\prime})=w(U)=w(V)$, and therefore $V\simeq U\oplus V^{\prime}$, as claimed. Since $\mathrm{T}(W)=U$, we have $V\simeq{\mathrm{T}}(W)\oplus V^{\prime}.$
By construction, since $X=\mathrm{T}(Y)$ has signature $(1,d-1)$, there is an embedding $\sigma:E\to\mathbf{R}$ with $Y_{\sigma}$ positive. Since $U_{2}=\mathrm{T}(W_{2})$ is negative definite, it follows that $(Y\oplus Y\oplus W_{2})_{\sigma}$ has signature $(2,m-2)$. ∎
@@
$V\simeq{\mathrm{T}}(W)\oplus V^{\prime}\leavevmode\nobreak\ .$
Moreover $W$ can be chosen in such a way that there is an embedding $\sigma:E\to\mathbf{R}$ with $W_{\sigma}$ of signature $(2,m-2)$ while all other embeddings $\tau:E\to\mathbf{R}$ have $W_{\tau}$ negative-definite.
@@
Let $h>0$ be such that $-2kh\in\mathrm{N}_{E/{\bf Q}}(E^{\times})$, and set $V^{\prime}=\langle h\rangle$. Then there exists a quadratic form $W$ over $E$ such that
$V\simeq{\mathrm{T}}(W)\oplus V^{\prime}\leavevmode\nobreak\ .$
Moreover $W$ can be chosen in such a way that there is an embedding $\sigma:E\to\mathbf{R}$ with $W_{\sigma}$ of signature $(2,m-2)$.
Conversely, if $V\simeq{\mathrm{T}}(W)\oplus V^{\prime}$ for some quadratic form $W$ over $E$ and $V^{\prime}=\langle h\rangle$, then $-2kh\in\mathrm{N}_{E/{\bf Q}}(E^{\times})$.
Proof. Set $U_{1}=H^{2}$. There exists a quadratic form $W_{1}$ over $E$ such that $\mathrm{T}(W_{1})=U_{1}$ and an embedding $\sigma:E\to\mathbf{R}$ with $(W_{1})_{\sigma}$ of signature $(2,2)$ (see [vGS 23], 3.11). Set $U^{\prime}=H\oplus\langle-2k\rangle$ or $H\oplus I_{16}\oplus\langle-2k\rangle$. The quadratic form $H$ represents all non-zero rational numbers, hence there exists $x\in U^{\prime}$ such that $q(x,x)=h$. Let $U_{2}$ be such that $U^{\prime}\simeq U_{2}\oplus\langle h\rangle$. By [Kr 92], Proposition 6, there exists a form $W_{2}$ such that $\mathrm{T}(W_{2})=U_{2}$. Set $W=W_{1}\oplus W_{2}$. and $U=U_{1}\oplus U_{2}$. By construction, we have $V\simeq{\mathrm{T}}(W)\oplus V^{\prime}$, and the signature of $W_{\sigma}$ is $(2,m-2)$.)
Conversely, assume that $V\simeq{\mathrm{T}}(W)\oplus V^{\prime}$. This implies that ${\rm det}({\mathrm{T}}(W))={\rm det}(V){\rm det}(V^{\prime})=2kh$. Since $E$ is a quadratic field, $-\Delta_{E}\in{\mathrm{N}}_{E/{\bf Q}}(E^{\times})$, and Lemma 5.1 (ii) implies that $-{\rm det}({\mathrm{T}}(W))\in{\mathrm{N}}_{E/{\bf Q}}(E^{\times})$, hence $-2kh\in{\mathrm{N}}_{E/{\bf Q}}(E^{\times})$, as claimed. ∎
@@ 8. Witt groups
The aim of this section is to recall some notions and results concerning Witt rings of quadratic forms; see for instance [MH 73] for details. Let $F$ be a field of characteristic $2$. We denote by ${\rm Witt}(F)$ the Witt ring of $F$, and by $I(F)$ the fundamental ideal of ${\rm Witt}(F)$, i.e. the ideal of the even dimensional quadratic forms.
@@
Suppose that $E$ is a totally real number field, and let $U\in I({\mathbf{Q}})$ be a torsion form. There exists $W\in I(E)$ such that $T(W)\simeq U$ if and only if ${\rm disc}(U)\in\Lambda^{+}_{E/F}$.
Proof. By a result of Krüskemper [Kr 92], Lemma 7, such a $W\in I(E)$ exists if and only if ${\rm N}_{E/{\bf Q}}(a)={\rm disc}(U)$ in ${\mathbf{Q}}^{\times}/{{\mathbf{Q}}}^{\times 2}$ for some $a\in E^{\times}$ that is a sum of squares in $E$. Since $E$ is totally real, an element $a\in E^{\times}$ is a sum of squares if and only if it is totally positive (see for instance [L 73], Chapter VII, Theorem 1.12). ∎
@@
Suppose that $E$ is a totally real number field, and let $V\in{\rm Witt}(E)$. Then $E$ is a torsion element of the group ${\rm Witt}(E)$ if and only if ${\rm ind}_{\sigma}(V)=0$ for all $\sigma\in\Sigma_{E}$.
@@ 9. Characterization
Let $U$ be a quadratic form over ${\mathbf{Q}}$ of dimension $r$ and signature $(2,r-2)$. Let $E$ be a number field of degree $d$, and let $m$ be an integer such that $r=md$; assume that either $E$ is CM and $m\geqslant 1$ or $E$ is totally real and $m\geqslant 3$.
Question 9.1.
@@
$\bullet$ $\mathrm{T}(W)\simeq U$;
$\bullet$ If $E$ is totally real, is there a $\sigma\in\Sigma_{E}$ such that the signature of $W_{\sigma}$ is $(2,m-2)$
@@
(i) ${\rm dim}_{{\mathbf{Q}}}(U)=m[E:{\mathbf{Q}}]$.
(ii) ${\rm disc}(U)=\Delta_{E}^{m}$;
(iii) If $p\in S_{E}$, then $U\otimes_{{\mathbf{Q}}}{{\mathbf{Q}}}_{p}$ is isomorphic to an orthogonal sum of hyperbolic planes.
(iv) The signature of $U$ is of the form $(2a,2b)$ for some integers $a,b\geqslant 0$.
Proof. This follows from Theorem 5.5. Indeed, ${\rm dim}(U)=2d_{0}m$, hence ${\rm disc}(U)=(-1)^{d_{0}m}{\rm det}(V)$; this implies that condition (ii) above and condition (ii) of Theorem 5.5 are equivalent. ∎
These conditions are rather restrictive; as we will see, the case of totally real fields is quite different. We start with the case where $d$ is odd, and apply a method of Krüskemper, [Kr 92].
@@
Moreover $W$ can be chosen in such a way that there is an embedding $\sigma:E\to\mathbf{R}$ with $W_{\sigma}$ of signature $(2,m-2)$.
@@
We follow the proof of Krüskemper in [Kr 92], Proof of Theorem 3 (b), page 115. Let us consider the Witt class $X=\mathrm{T}(W^{\prime})-U$ in ${\rm Witt}({\mathbf{Q}})$, and note that $X\in I({\mathbf{Q}})$. By [Kr 92], Corollary of Lemma 7, page 114, there exists $Y\in I(E)$ such that $\mathrm{T}(Y)=X$. Let $W$ be a quadratic form of dimension $m$ over $E$ representing the Witt class $Y-W^{\prime}$. This is possible since $m\geqslant 3$, see for instance [O’M 73], 72:1 and 63:23 (these results generalize Theorem 1.1 to number fields). Then $\mathrm{T}(W)=U$ in ${\rm Witt}({\mathbf{Q}})$, and since ${\rm dim}(\mathrm{T}(W))={\rm dim}(U)$, we have $\mathrm{T}(W)\simeq U$. ∎
@@
Moreover, assume that there is an embedding $\sigma:E\to\mathbf{R}$ with $W_{\sigma}$ of signature $(2,m-2)$.
@@
Moreover $W$ can be chosen in such a way that there is an embedding $\sigma:E\to\mathbf{R}$ with $W_{\sigma}$ of signature $(2,m-2)$.
@@
The element $\alpha\beta$ is totally positive, and $\Delta_{E}^{2m}$ is also the norm of a totally positive element; hence by Theorem 8.1 there exists a torsion form $Y\in I(E)$ such that $\mathrm{T}(Y)=X$. Let $W$ be a quadratic form of dimension $m$ over $E$ representing the Witt class $Y-W^{\prime}$. Then $\mathrm{T}(W)=U$ in ${\rm Witt}({\mathbf{Q}})$, and since ${\rm dim}(\mathrm{T}(W))={\rm dim}(U)$, we have $\mathrm{T}(W)\simeq U$. ∎
@@
Let $E$ be a totally real field. If $E/{\bf Q}$ is a Galois extension, then $\Lambda^{+}_{E/{{\mathbf{Q}}}}=\Lambda_{E/{{\mathbf{Q}}}}$.
@@
Indeed, $d$ is a sum of two squares $\iff$ $d\in{\rm N}_{{\mathbf{Q}}(i)/{{\mathbf{Q}}}}({{\mathbf{Q}}(i)})$ $\iff$ $(-1,d)=0$ $\iff$ $-1\in{\rm N}_{E/{{\mathbf{Q}}}}(E)$ $\iff$ $d\in{\rm N}_{E/{{\mathbf{Q}}}}(E)$ $\iff$ $\Delta_{E}\in{\rm N}_{E/{{\mathbf{Q}}}}(E)$ $\iff$ ${\rm det}(U)\in\Lambda^{+}_{E/{{\mathbf{Q}}}}\Delta_{E}^{m}=\Lambda_{E/{{\mathbf{Q}}}}\Delta_{E}$; the last equality follows from Lemma 9.6. By Theorem 9.5, we have ${\rm det}(U)\in\Lambda^{+}_{E/{{\mathbf{Q}}}}\Delta_{E}^{m}$ $\iff$ there exists an $m$-dimensional quadratic form $W$ over $E$ such that $\mathrm{T}(W)\simeq U$.

### v3

@@ 8. Totally real fields
If $E$ is a totally real field, we denote by $\Sigma_{E}$ be the set of real embeddings of $E$; we have $E\otimes_{{\mathbf{Q}}}\mathbf{R}=\prod_{\sigma\in\Sigma_{E}}E_{\sigma}$, with $E_{\sigma}={\bf R}$ for all $\sigma\in\Sigma_{E}$. If $W$ is a quadratic form over $E$, then $W\otimes_{{\mathbf{Q}}}\mathbf{R}$ decomposes as an orthogonal sum $W\otimes_{{\mathbf{Q}}}\mathbf{R}=\bigoplus_{\sigma\in\Sigma_{E}}W_{\sigma}$; each of the $W_{\sigma}$ is a quadratic form over $\mathbf{R}$.
@@
  • $\bullet$
    $md<{\rm dim}(V)-2$,
  • $\bullet$
    $md={\rm dim}(V)-2$, and moreover for $p=2$ and for all prime numbers $p$ dividing ${\rm N}_{E/{\mathbf{Q}}}({\rm det}(W))$, we have ${\rm det}(\mathrm{T}(W))\not=-{\rm det}(V)$ in ${\mathbf{Q}}_{p}^{\times}/{\mathbf{Q}}_{p}^{\times 2}$.
@@
$V\simeq{\mathrm{T}}(W)\oplus V^{\prime}\penalty 10000\ .$
Moreover $W$ can be chosen in such a way that there is an embedding $\sigma:E\to\mathbf{R}$ with $W_{\sigma}$ of signature (2,m2)(2,m-2).
Proof. Let $X$ be a quadratic form over $\mathbf{Q}$} of signature (1,d1)(1,d-1) such that ${\rm det}(X)=1$ and $w(X)=1$ at the prime $2$ and at infinity, and 0 elsewhere; this is possible by Theorem 2.1, (i). By a result of Krüskemper [Kr 92] Theorem A there exists a 1-dimensional quadratic form $Y$ over $E$ of such that $\mathrm{T}(Y)\simeq X$; indeed, the hypotheses of this result are satisfied, since $w(X)=0$ at all non-dyadic primes. Set $U_{1}=X\oplus X$ and $W_{1}=Y\oplus Y$; we have $\mathrm{T}(W_{1})\simeq U_{1}$. The quadratic form $U_{1}$ has signature (2,2d2)(2,2d-2), trivial determinant and Hasse invariant.
Let $U_{2}$ be a negative definite quadratic form over $\mathbf{Q}$} of dimension (m2)d(m-2)d, determinant h-h and such that $w(U_{2})=w(V)$; such a quadratic form exists by Theorem 2.1 (ii). There exists a quadratic form $W_{2}$ over $E$ of such that $\mathrm{T}(W_{2})\simeq U_{2}$; this follows from [Kr 92], Theorem A if $m=3$, and from [Kr 92], Theorem 3 b) if $m=7$.
Set $W=W_{1}\oplus W_{2}$ and $U=U_{1}\oplus U_{2}$. Note that $U$ is a quadratic form over $\mathbf{Q}$} of signature $(2,19)$ with ${\rm det}(U)=-h$ and $w(U)=w(V)$. We have $V^{\prime}=\langle-{\rm det}(U)\rangle$; we claim that $V\simeq U\oplus V^{\prime}$. Indeed, these forms have the same dimension, determinant and signature. We have $w(U\oplus V^{\prime})=w(U)+w(V^{\prime})+({\rm det}(U),{\rm det}(V^{\prime}))$. Since ${\rm det}(V^{\prime})=-{\rm det}(U)$, we have $({\rm det}(U),{\rm det}(V^{\prime}))=0$. Moreover, ${\rm dim}(V^{\prime})=1$, hence $w(V^{\prime})=0$. This implies that $w(U\oplus V^{\prime})=w(U)=w(V)$, and therefore $V\simeq U\oplus V^{\prime}$, as claimed. Since $\mathrm{T}(W)=U$, we have $V\simeq{\mathrm{T}}(W)\oplus V^{\prime}.$
By construction, since $X=\mathrm{T}(Y)$ has signature (1,d1)(1,d-1), there is an embedding $\sigma:E\to\mathbf{R}$ with $Y_{\sigma}$ positive. Since $U_{2}=\mathrm{T}(W_{2})$ is negative definite, it follows that $(Y\oplus Y\oplus W_{2})_{\sigma}$ has signature (2,m2)(2,m-2). ∎
@@
$V\simeq{\mathrm{T}}(W)\oplus V^{\prime}\penalty 10000\ .$
Moreover $W$ can be chosen in such a way that there is an embedding $\sigma:E\to\mathbf{R}$ with $W_{\sigma}$ of signature (2,m2)(2,m-2) while all other embeddings $\tau:E\to\mathbf{R}$ have $W_{\tau}$ negative-definite.
@@
Let $h>0$ be such that $2kh\in\mathrm{N}_{E/{\bf Q}}(E^{\times})$, and set $V^{\prime}=\langle h\rangle$. Then there exists a quadratic form $W$ over $E$ such that
$V\simeq{\mathrm{T}}(W)\oplus V^{\prime}\penalty 10000\ .$
Moreover $W$ can be chosen in such a way that there is an embedding $\sigma:E\to\mathbf{R}$ with $W_{\sigma}$ of signature (2,m2)(2,m-2).
Conversely, if $V\simeq{\mathrm{T}}(W)\oplus V^{\prime}$ for some quadratic form $W$ over $E$ and $V^{\prime}=\langle h\rangle$, then $2kh\in\mathrm{N}_{E/{\bf Q}}(E^{\times})$.
Proof. Set $U_{1}=H^{2}$. There exists a quadratic form $W_{1}$ over $E$ such that $\mathrm{T}(W_{1})=U_{1}$ and an embedding $\sigma:E\to\mathbf{R}$ with $(W_{1})_{\sigma}$ of signature $(2,2)$ (see [vGS 25], 3.11). Set $U^{\prime}=H\oplus\langle-2k\rangle$ or $H\oplus I_{16}\oplus\langle-2k\rangle$. The quadratic form $H$ represents all non-zero rational numbers, hence there exists $x\in U^{\prime}$ such that $q(x,x)=h$. Let $U_{2}$ be such that $U^{\prime}\simeq U_{2}\oplus\langle h\rangle$. By [Kr 92], Proposition 6, there exists a form $W_{2}$ such that $\mathrm{T}(W_{2})=U_{2}$. Set $W=W_{1}\oplus W_{2}$. and $U=U_{1}\oplus U_{2}$. By construction, we have $V\simeq{\mathrm{T}}(W)\oplus V^{\prime}$, and the signature of $W_{\sigma}$ is (2,m2)(2,m-2).
Conversely, assume that $V\simeq{\mathrm{T}}(W)\oplus V^{\prime}$. This implies that ${\rm det}({\mathrm{T}}(W))={\rm det}(V){\rm det}(V^{\prime})=2kh$. Since $E$ is a quadratic field, $\Delta_{E}\in{\mathrm{N}}_{E/{\bf Q}}(E^{\times})$, and Lemma 6.1 (ii) implies that ${\rm det}({\mathrm{T}}(W))\in{\mathrm{N}}_{E/{\bf Q}}(E^{\times})$, hence $2kh\in{\mathrm{N}}_{E/{\bf Q}}(E^{\times})$, as claimed. ∎
@@ 9. Witt groups
The aim of this section is to recall some notions and results concerning Witt rings of quadratic forms; see for instance [L 05] or [MH 73] for details. Let $F$ be a field of characteristic $\not=2$. We denote by ${\rm Witt}(F)$ the Witt ring of $F$, and by $I(F)$ the fundamental ideal of ${\rm Witt}(F)$, i.e. the ideal of the even dimensional quadratic forms.
The previously defined invariants, such as dimension, determinant, signature and Hasse invariant, do not vanish on hyperbolic forms, hence they are not well-defined on ${\rm Witt}(F)$. Therefore, one has to modify these invariants. First of all, the dimension is only defined modulo $2$: in the sequel, the dimension of a Witt class will be understood as an element of ${\mathbf{Z}}/2{\mathbf{Z}}$. Instead of the determinant, we consider the discriminant, defined as follows.
@@
Next, we modify the Hasse invariant, as in [L 05, Chapter V, Propositions 3.19, 3.20], in order to obtain an invariant
(9.1) $c:{\rm Witt}(F)\to\mathrm{Br}_{2}(F),$
called the Witt invariant. The relationship with the Hasse invariant can be expressed in a (rather complicated) formula, see [L 05, Chapter V, Proposition 3.20].
@@
Example 9.1.
Let $E={\mathbf{Q}}$. The form $V=\langle 1,-5\rangle$ does not represent 0 in $\rm{Witt}({\mathbf{Q}})$, but it is a torsion form, of order $2$. Indeed, the form $V\oplus V$ has dimension 4, signature $(2,2)$, determinant $1$ (mod squares), while its Hasse invariant is (5,5)=(1,5)(-5,-5)=(-1,-5). This is non-zero at 2 and over the real numbers, and zero everywhere else locally, hence it is equal to the Hasse invariant of the hyperbolic form $H^{2}$. This implies that all the invariants of this form coincide with those of $H^{2}$, hence it is isomorphic to $H^{2}$.
@@
Suppose that $E$ is a totally real number field, and let $U\in I({\mathbf{Q}})$ be a torsion form. There exists a torsion form $W\in I(E)$ such that $T(W)\simeq U$ if and only if ${\rm disc}(U)\in\Lambda^{+}_{E/F}$.
Proof. By a result of Krüskemper [Kr 92], Lemma 7, such a $W\in I(E)$ exists if and only if ${\rm N}_{E/{\bf Q}}(a)={\rm disc}(U)$ in ${\mathbf{Q}}^{\times}/{{\mathbf{Q}}}^{\times 2}$ for some $a\in E^{\times}$ that is a sum of squares in $E$. Since $E$ is totally real, an element $a\in E^{\times}$ is a sum of squares if and only if it is totally positive (see for instance [L 05], Chapter VII, Theorem 1.12). ∎
@@
Suppose that $E$ is a totally real number field, and let $V\in{\rm Witt}(E)$. Then $V$ is a torsion element of the group ${\rm Witt}(E)$ if and only if ${\rm ind}_{\sigma}(V)=0$ for all $\sigma\in\Sigma_{E}$.
@@ 10. Characterization
Let $U$ be a quadratic form over $\mathbf{Q}$} of dimension $r$ and signature (2,r2)(2,r-2). Let $E$ be a number field of degree $d$, and let $m$ be an integer such that $r=md$; assume that either $E$ is CM and $m\geqslant 1$ or $E$ is totally real and $m\geqslant 3$.
Question 10.1.
@@
  • $\bullet$
    $\mathrm{T}(W)\simeq U$;
  • $\bullet$
    If $E$ is totally real, is there a $\sigma\in\Sigma_{E}$ such that the signature of $W_{\sigma}$ is (2,m2)(2,m-2)?
@@
  1. (1)
    ${\rm dim}_{{\mathbf{Q}}}(U)=m[E:{\mathbf{Q}}]$.
  2. (2)
    ${\rm disc}(U)=\Delta_{E}^{m}$;
  3. (3)
    If $p\in S_{E}$, then $U\otimes_{{\mathbf{Q}}}{{\mathbf{Q}}}_{p}$ is isomorphic to an orthogonal sum of hyperbolic planes.
  4. (4)
Proof. This follows from Theorem 6.5. Indeed, ${\rm dim}(U)=2d_{0}m$, hence ${\rm disc}(U)=(-1)^{d_{0}m}{\rm det}(U)$; this implies that condition (ii) above and condition (ii) of Theorem 6.5 are equivalent. ∎
These conditions are rather restrictive; as we will see, the case of totally real fields is quite different. Before we proceed, we need another auxiliary result.
Lemma 10.3.
Let $E$ be a number field, let $W\in{\rm Witt}(E)$ and let $m\geq 3$ be an integer. Suppose that $\dim(W)\equiv m\mod 2$, and that the indices of $W$ have absolute value $|{\rm ind}_{\sigma}(W)|\leq m$ at all the real places $\sigma$ of $E$. Then there exists a quadratic form $V$ over $E$ of dimension $m$ representing the Witt class $W$.
Proof. Let $c\in\mathrm{Br}_{2}E$ be the Witt invariant of $W$ from Section 9, and let $w\in\mathrm{Br}_{2}E$ correspond to $c$ as in (9.1). Since $m\geq 3$ and the indices of $W$ are $\leq m$ at all the real places of E, there exists a quadratic form $V$ of dimension $m$ over $E$ of Hasse invariant $w$, with the same indices and discriminant as $W$; this follows from [O’M 73, 72:1 and 63:23] (these results generalize Theorem 2.1 to number fields). The form $V$ represents the Witt class $W$. ∎
Concentrating on totally real fields, we start with the case where $d$ is odd, and apply a method of Krüskemper, [Kr 92].
@@
Moreover $W$ can be chosen in such a way that there is an embedding $\sigma:E\to\mathbf{R}$ with $W_{\sigma}$ of signature (2,m2)(2,m-2).
@@
We follow the proof of Krüskemper in [Kr 92], Proof of Theorem 3 (b), page 115. Let us consider the Witt class $X=\mathrm{T}(W^{\prime})-U$ in ${\rm Witt}({\mathbf{Q}})$, and note that $X\in I({\mathbf{Q}})$. By [Kr 92], Corollary of Lemma 7, page 114, there exists a torsion class $Y\in I(E)$ such that $\mathrm{T}(Y)=X$. (Note that applying this corollary requires the degree $d$ to be odd.) Since $Y$ is torsion, its indices are 0 at all the real places of $E$ (see Theorem 9.3). This implies that the indices of $Y-W^{\prime}$ are $\leq m$ at all the real places of $E$. Note that the dimension of $Y$ is even, since $Y\in I(E)$, and that $\dim(W^{\prime})=m$, hence the dimension of the Witt class $Y-W^{\prime}$ is congruent to $m$ modulo $2$. Since moreover $m\geq 3$, we can apply Lemma 10.3 and conclude that there exists a quadratic form $W$ of dimension $m$ over $E$ representing the Witt class $Y-W^{\prime}$. Then $\mathrm{T}(W)=U$ in ${\rm Witt}({\mathbf{Q}})$, and since ${\rm dim}(\mathrm{T}(W))={\rm dim}(U)$, we have $\mathrm{T}(W)\simeq U$. ∎
@@
Moreover, assume that there is an embedding $\sigma:E\to\mathbf{R}$ with $W_{\sigma}$ of signature (2,m2)(2,m-2).
@@
Moreover $W$ can be chosen in such a way that there is an embedding $\sigma:E\to\mathbf{R}$ with $W_{\sigma}$ of signature (2,m2)(2,m-2).
@@
The element $\alpha\beta$ is totally positive, and $\Delta_{E}^{2m}$ is a square; hence the discriminant of $X$ belongs to $\Lambda^{+}{E/Q}$. Therefore, by Theorem 9.2 there exists a torsion form $Y\in I(E)$ such that $\mathrm{T}(Y)=X$. Let $W$ be a quadratic form of dimension $m$ over $E$ representing the Witt class $Y-W^{\prime}$; this exists by Lemma 10.3 as in the proof of Theorem 10.4 because $m>2$. Then $\mathrm{T}(W)=U$ in ${\rm Witt}({\mathbf{Q}})$, and since ${\rm dim}(\mathrm{T}(W))={\rm dim}(U)$, we have $\mathrm{T}(W)\simeq U$. ∎
@@
Let $E$ be a totally real field. If $E/{\bf Q}$ is a Galois extension and $a\in\Lambda_{E/{{\mathbf{Q}}}}$ satisfies $a>0$, then $a\in\Lambda^{+}_{E/{{\mathbf{Q}}}}$.
@@
Indeed, $d$ is a sum of two squares $\iff$ $\Delta_{E}\in{\rm N}_{E/{{\mathbf{Q}}}}(E)$ $\iff$ ${\rm det}(U)\in\Lambda_{E/{{\mathbf{Q}}}}\Delta_{E}\iff\det(U)\in\Lambda^{+}_{E/{{\mathbf{Q}}}}\Delta_{E}^{m}$; here the first equivalence was covered in Example 1.5, the second equivalence is obvious, and the last equivalence follows from Lemma 10.7. By Theorem 10.6, we have ${\rm det}(U)\in\Lambda^{+}_{E/{{\mathbf{Q}}}}\Delta_{E}^{m}$ $\iff$ there exists an $m$-dimensional quadratic form $W$ over $E$ such that $\mathrm{T}(W)\simeq U$.